Fermi National Laboratory
 

Operation Summary for October 8, 2010 - October 11, 2010

 

The Integrated Luminosity for the period from 10/4/10 to 10/11/10 was 48.65 pb-1 . NuMI
reported receiving no protons on target during this same period due to target work.

Friday Oct. 8
The day shift began with Operations monitoring the Tevatron (TeV) store 8152 (122.74E30) and the Recycler stash (183.1E10), with theAntiproton Source (Pbar) stacking (23.2E10), with the Linac, Booster, and theMain Injector(MI) operating normally, with NuMI off for target work, with MiniBooNE taking beam, and with theCollider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding store.

During the day shift, Operations terminated store 8152 at 12:16 PM. Operators prepared D0 for an access, which concluded at 1:13 PM.

During the evening shift, Operations put the TeV into HEP shot setup, but the TeV quenched at 4:22 PM due to tuning implemented between stores. The TeV sector A4 wet engine failed immediately following the quench, which required cryo system technicians to begin a full engine overhaul and other work.

Saturday Oct. 9
The midnight shift began with the TeV recovering from a quench, with Operations monitoring the stash (240.8E10), with Pbar stacking (13.4E10), and with MiniBooNE taking beam.

During the midnight shift, cryo system techs continued working on the A4 dry engine and cryo system.

During the day shift, cryo system techs completed their dry engine work at 1:13 PM. Operators began the TeV turn-on and checkout and then put the TeV into shot setup.

During the evening shift, Operations established store 8159 at 5:04 PM with an average initial luminosity (AIL) of 296.47E30.

Sunday Oct. 10
The midnight shift began with Operations monitoring store 8159 and the stash, with Pbar stacking, and with MiniBooNE taking beam.

During the midnight shift, Operations terminated store 8159 at 5:13 AM, put the TeV into shot setup, and established store 8160 at 6:30 AM with an AIL of 320.5E30.

During the day shift, Operations terminated store 8160 at 3:39 PM and put the TeV into shot setup.

During the evening shift, Operations established store 8161 at 5 PM with an AIL of 327.49E30.

Monday Oct. 11
The midnight shift began with Operations monitoring store 8161 and the stash, with Pbar stacking, and with MiniBooNE taking beam.

During the midnight shift, Operations terminated store 8161 at 6:36 AM and put the TeV into shot setup.

The Plans for Monday
The plans for today are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/11/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory